OpenStudy (anonymous):

What is the 34th term of the following sequence? 2,7,12,...

OpenStudy (lgbasallote):

hint: first term = 2 total number of terms = 34 common difference = 5 what you're looking for = 34th term use \[a_n = a_ 1 + (n-1)d\] dooes that help?

OpenStudy (anonymous):

No it's a new subject. Can you do it for me so I'll understand how to do my others?

OpenStudy (anonymous):

Or show me the steps

OpenStudy (lgbasallote):

how about an example 2,8,14, 20.... i would like to find the 30th term now i use \[a_n = a_1 + (n-1)d\] a_n means the last term a_1 means the first term n is the number of terms d is the common difference a_n is what im looking for in this case a_1 is 2 because that's the first term in my given n is 30 because im looking for the 30th term obviously the number of terms would be 30 d is 6 because if you notice im adding 6 to the current term to get the next term so i use the formula.. \[a_n = a_1 + (n-1)d\] \[\implies a_{32} = 2 + (30-1)6\] \[\implies a_{32} = 2 + (29)(6)\] \[\implies a_{32} = 2 + 174\] \[\implies a_{32} = 176\] does that help?

OpenStudy (anonymous):

I'm going to try to fit this into that formula hold on..

OpenStudy (lgbasallote):

sure

OpenStudy (lgbasallote):

just tag me if you need me

OpenStudy (anonymous):

@lgbasallote Where did you get 32 from for the base?

OpenStudy (lgbasallote):

uhh lol that should be 30...that mans 30th term =))

OpenStudy (anonymous):

Oh that makes more sense

OpenStudy (anonymous):

change it before you confuse other people lol

OpenStudy (lgbasallote):

sorry bout that

OpenStudy (anonymous):

I got a_34= 167 Thank you

OpenStudy (lgbasallote):

welcome